what is the value of x if 1/4 x-1/8 x=3- 1/16 x
Answer by mananth(16946)

1/4 x-1/8 x=3- 1/16 x
1/4x - 1/8x +1/16x =3
LCM = 16x
4-2+1 /16x= 3
3= 48x
x=3/48
x=1/16